OpenStudy (anonymous):

Find the sum of the geometric sequence. 1 divided by 3, 2 divided by 3, 4 divided by 3, 8 divided by 3, 16 divided by 3 31 divided by 3 10 2 31 divided by 15

OpenStudy (anonymous):

@hoblos @eliassaab @ganeshie8

OpenStudy (anonymous):

will give best answer & medal!

OpenStudy (anonymous):

Here's the geometric sequence: 1/3 , 2/3 , 4/3 , 8/3, 16/3

ganeshie8 (ganeshie8):

\(\large \frac{1}{3}, \frac{2}{3}, \frac{4}{3}, \frac{8}{3}, \frac{16}{3}\)

OpenStudy (anonymous):

the answer choices: A) 31/3 B) 10 C) 2 D) 31/15

ganeshie8 (ganeshie8):

either add them up manually or use partial series formula

OpenStudy (anonymous):

kk! when I add them up manually i got 10 (:

OpenStudy (anonymous):

is that the correct answer?

ganeshie8 (ganeshie8):

\(\large \frac{1}{3} + \frac{2}{3} + \frac{4}{3} + \frac{8}{3} + \frac{16}{3}\) \(\large \frac{1}{3}\left(1 + 2 + 4 + 8 + 16\right)\)

ganeshie8 (ganeshie8):

try again

OpenStudy (anonymous):

I still got 10 lol :(

ganeshie8 (ganeshie8):

\(\large \frac{1}{3} + \frac{2}{3} + \frac{4}{3} + \frac{8}{3} + \frac{16}{3}\) \(\large \frac{1}{3}\left(1 + 2 + 4 + 8 + 16\right)\) \(\large \frac{1}{3}\left(31\right)\)

OpenStudy (anonymous):

Yes that's equal to 10.323

ganeshie8 (ganeshie8):

10.323 is NOT equal to 10

ganeshie8 (ganeshie8):

look at ur options, u have 31/3 also right ?

OpenStudy (anonymous):

Sorry my teacher makes us round our answers!

OpenStudy (anonymous):

Ohhhhhhhhhhhhh good catch!

OpenStudy (anonymous):

thank u SOO much!

OpenStudy (anonymous):

:)

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n. 8 + 16 + 24 + ... + 8n = 4n(n + 1)

OpenStudy (anonymous):

@ganeshie8 can u help me with this one?

OpenStudy (anonymous):

since u seem to definatley know what ur doing! lol

OpenStudy (anonymous):

:)

ganeshie8 (ganeshie8):

3 steps : 1) show that the formula works for \(n = 1\) 2) assume the formula works for some \(n=k\) 3) show that it works for \(n=k+1\)

ganeshie8 (ganeshie8):

^^thats the basic setup for any induction proof

OpenStudy (anonymous):

Can u show me the steps written out? I've never done this before

ganeshie8 (ganeshie8):

\(8 + 16 + 24 + ... + 8n = 4n(n + 1)\) step 1 : lets see if the formula works for \(n=1\) case

ganeshie8 (ganeshie8):

plugin \(n=1\), on right hand side formula wat do u get ?

OpenStudy (anonymous):

4(1)(2) 8 :)

ganeshie8 (ganeshie8):

so, n =1 case lies in our solution set for formula. move to step 2

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

that's the trickiest part

ganeshie8 (ganeshie8):

\(8 + 16 + 24 + ... + 8n = 4n(n + 1)\) step 1 : lets see if the formula works for \(n=1\) case \(4*1(1+1) = 4(2) = 8\) So, the formula is \(true\) for sum of \(1\) terms. step 2 : assume that the formula for works for \(n=k\) then : \(8 + 16 + 24 + ... + 8k = 4k(k + 1)\)

ganeshie8 (ganeshie8):

actually, step2 is the easiest step in any induction proof :- u just need to assume that the formula works for n = k, and replace n wid k in the formula... thats all. move to step 3

OpenStudy (anonymous):

oh ok!

ganeshie8 (ganeshie8):

step 3 can be a bit tricky, as u need to show that the formula works for n = k+1, whenever it worked for n = k

ganeshie8 (ganeshie8):

its not that long process, its just 3 steps :)

ganeshie8 (ganeshie8):

once u do 1-2 problems, u will start to like these proofs

OpenStudy (anonymous):

got it :)

ganeshie8 (ganeshie8):

\(8 + 16 + 24 + ... + 8n = 4n(n + 1)\) step 1 : lets see if the formula works for \(n=1\) case \(4*1(1+1) = 4(2) = 8\) So, the formula is \(true\) for sum of \(1\) terms. step 2 : assume that the formula for works for \(n=k\) then : \(8 + 16 + 24 + ... + 8k = 4k(k + 1)\) step 3 : sum of \(k+1\) terms is \(8 + 16 + 24 + ... + 8(k+1)\)

ganeshie8 (ganeshie8):

you need to somehow show that it equals : \(4(\color{red}{k+1})(\color{red}{k+1}+1)\)

ganeshie8 (ganeshie8):

any idea, why we need to show that ?

OpenStudy (anonymous):

hmm, its to prove the formula true for all positive integeres

OpenStudy (anonymous):

I'm not sure

ganeshie8 (ganeshie8):

you wanted to prove : \(8 + 16 + 24 + ... + 8n = 4n(n + 1)\)

ganeshie8 (ganeshie8):

plugin n = (k+1) wat do u get ?

OpenStudy (anonymous):

The answer u posted in red above!

ganeshie8 (ganeshie8):

where ever u see "n", just blindly put "k+1"

OpenStudy (anonymous):

4(k+1)(k+1+1) :)

OpenStudy (anonymous):

will do!

ganeshie8 (ganeshie8):

yup!

ganeshie8 (ganeshie8):

so, \(8 + 16 + 24 + ... + 8(k+1) = 4(k+1)((k+1) + 1) \) is the sum formula for k+1 terms

ganeshie8 (ganeshie8):

lets go back to the proof

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

btw your an awesome teacher :)

ganeshie8 (ganeshie8):

\(8 + 16 + 24 + ... + 8n = 4n(n + 1)\) step 1 : lets see if the formula works for \(n=1\) case \(4*1(1+1) = 4(2) = 8\) So, the formula is \(true\) for sum of \(1\) terms. step 2 : assume that the formula for works for \(n=k\) then : \(8 + 16 + 24 + ... + 8k = 4k(k + 1)\) step 3 : sum of \(k+1\) terms is \(8+16+24+...+8(k+1) = 8+16+24+...+8k+8 \) from step2 of induction hypothesis, this equals : \( 4k(k+1)+8 \)

ganeshie8 (ganeshie8):

we're almost done, just make sure you're at peace wid above^

ganeshie8 (ganeshie8):

next u need to show \(4k(k+1) + 8\) is same as \(4(k+1)(k+1 + 1)\)

OpenStudy (anonymous):

yes definatley at pace :) sorry if im slow im taking notes but i only have 10 mins left on the computer

ganeshie8 (ganeshie8):

\(4k(k+1) + 8\) \(4k^2+4k + 8\) \(4(k^2+k + 2)\)

ganeshie8 (ganeshie8):

factor the thing inside parenthesis, wat do u get ?

OpenStudy (anonymous):

is it factorable?

OpenStudy (anonymous):

Oh you mean factor it out 4k^2 + 4k +8

ganeshie8 (ganeshie8):

yup, factor below \(k^2 + k + 2\)

OpenStudy (anonymous):

Idk how to do that sorry :(

OpenStudy (anonymous):

You can do it really quick and I will search how to do it later due to the low time amount

ganeshie8 (ganeshie8):

actually i made a mistake earlier

ganeshie8 (ganeshie8):

in step 3

OpenStudy (anonymous):

whats the mistake?

ganeshie8 (ganeshie8):

\(8 + 16 + 24 + ... + 8n = 4n(n + 1)\) step 1 : lets see if the formula works for \(n=1\) case \(4*1(1+1) = 4(2) = 8\) So, the formula is \(true\) for sum of \(1\) terms. step 2 : assume that the formula for works for \(n=k\) then : \(8 + 16 + 24 + ... + 8k = 4k(k + 1)\) step 3 : sum of \(k+1\) terms is \(8+16+24+...8k+8(k+1) \) from step2 of induction hypothesis, this equals : \( 8+16+24+...8k+8(k+1) = 4k(k+1)+8(k+1) \) \(\qquad ~~~~ = 4k^2 + 4k + 8k + 8\) \(\qquad ~~~~ = 4k^2 + 12k + 8\) \(\qquad ~~~~ = 4(k^2 + 3k + 2)\) \(\qquad ~~~~ = 4(k+1)(k+2)\) \(\qquad ~~~~ = 4(k+1)(k+1 + 1)\) So, the formula works for \(n = k+1\) also. Hence the formula works for all \(n \ge 1\)

OpenStudy (anonymous):

thank u SOOO much! :)

ganeshie8 (ganeshie8):

^^complete proof above, let me knw if smthng doesnt make sense..

OpenStudy (anonymous):

will do! it should since ur a professional at this :)

ganeshie8 (ganeshie8):

lol no, u wil become professional after doing 2-3 proofs, trust me u wil get addicted to induction proofs once u feel how powerful they're lol

OpenStudy (anonymous):

lol ok :) thank u so much for ur help, I wish i could give you best response more than once

ganeshie8 (ganeshie8):

:) you're wlcme !